OpenStudy (abbyselena2199):

At the beginning of every year, Molly deposits $200 in a savings account that offers an interest rate of 20%, compounded annually. The total amount that Molly will have in her account at the end of 3 years is?

OpenStudy (abbyselena2199):

Medal will be given

rebeccaxhawaii (rebeccaxhawaii):

WELCOME TO OS!!!

rebeccaxhawaii (rebeccaxhawaii):

do you know the fomula?

OpenStudy (abbyselena2199):

NO! You dont help me anyway @rebeccaxhawaii

rebeccaxhawaii (rebeccaxhawaii):

i am being tagged in a bajillion questions

OpenStudy (anonymous):

A = P (1+r)^t A = $200 (1 + .20)^3 A= $200 ( 1.20)^3 A= $200 (1.728) A= $345.6